Find all numbers for which the rational expression is undefined:
s^3-9s/s^2-4
Please explain the answer.
PLEASE HELP!!!

Answer by nerdybill(7384)

Find all numbers for which the rational expression is undefined:
s^3-9s/s^2-4
.
.
If you divide anything by a zero, it is considered "undefined". That is, when the denominator equals zero your rational expression will be undefined.
.
So, if we set the denominator (s^2-4) to zero and solve for s we'll find your answer:
s^2-4 = 0
s^2 = 4
s = sqrt(4)
s = {-2,2} (this is what they are looking for)